OpenStudy (anonymous):

Integration by parts. Ive been studying for my final exam tomorrow in calc and iv seen two formulas for integration by parts. The one my teacher gave me is int_{?}^{?}[F'(x) * G(x)dx = int_{?}^{?} [F(x) * G(x)]'dx - int_{?}^{?} F(x) * G'(x) dx. But iv also been watching some videos on youtube teaching this and most of them use the formula int_{?}^{?}udv = uv - int_{?}^{?} vdu. Are they both the same or do they have different uses?

OpenStudy (anonymous):

sry every time you see int_{?}^{?} it is supposed to be integral. I thought it would work

OpenStudy (anonymous):

They are the same, as \[ \int \frac{d[F(x) G(x)]}{dx}dx = F(x)G(x)\]which is the same as the uv part done in other videos.

OpenStudy (anonymous):

k thats awsome. Cause i really like the udv equation better

OpenStudy (anonymous):

Indeed, it's quite clearer :-)